OpenStudy (anonymous):

Divide and simplify. (-15) / 3/5=?

jhonyy9 (jhonyy9):

-15*(5/3) = -5*5= -25

OpenStudy (anonymous):

This is actually a bit confusing and depends on the order of operation. My first guess is -1 which processing in way like most computer algebra software. You may see it here http://www.wolframalpha.com/input/?i=%28-15%29%2F3%2F5